A body with moment of inertia I = 10 kgm² rotates uniformly with T = 2s. Find the moment of momentum of the body, relative to the axis of rotation.


1
Expert's answer
2021-03-23T11:12:15-0400

By definition, the moment of momentum is:


"L = I\omega"

where "I = 10kg\cdot m^2" is the moment of inertia, and "\omega" is the angular speed. In turn, the angular speed is:


"\omega = \dfrac{2\pi}{T}"

where "T = 2s" is the period of rotation. Thus, obtian:


"L = \dfrac{2\pi I}{T}\\
L = \dfrac{2\pi \cdot 10}{2} = 10\pi\approx 31.4\space \dfrac{kg\cdot m^2}{s}"

Answer. "31.4\space \dfrac{kg\cdot m^2}{s}".